# ArchCAD-400K: An Open Large-Scale Architectural CAD Dataset and New Baseline for Panoptic Symbol Spotting

### v2

@@ Abstract
Recognizing symbols in architectural CAD drawings is critical for various advanced engineering applications. In this paper, we propose a novel CAD data annotation engine that leverages intrinsic attributes from systematically archived CAD drawings to automatically generate high-quality annotations, thus significantly reducing manual labeling efforts. Utilizing this engine, we construct ArchCAD-400K, a large-scale CAD dataset consisting of 413,062 chunks from 5538 highly standardized drawings, making it over 26 times larger than the largest existing CAD dataset. ArchCAD-400K boasts an extended drawing diversity and broader categories, offering line-grained annotations. Furthermore, we present a new baseline model for panoptic symbol spotting, termed Dual-Pathway Symbol Spotter (DPSS). It incorporates an adaptive fusion module to enhance primitive features with complementary image features, achieving state-of-the-art performance and enhanced robustness. Extensive experiments validate the effectiveness of DPSS, demonstrating the value of ArchCAD-400K and its potential to drive innovation in architectural design and construction.
@@ 1 Introduction
For a long time, CAD drawings have served as the universal language of architectural design, enabling seamless communication among designers, engineers, and construction personnel through standardized graphical primitives such as arcs, circles, and polylines. Accurate perception of these primitives in 2D CAD drawings is essential for various downstream applications, including automated drawing review and 3D building information modeling (BIM) [1, 2, 3, 4], as this ability enables the optimization of CAD-based workflows while enhancing efficiency and precision in architectural design and construction processes[5, 6].
Building upon earlier research in symbol spotting [7, 8, 9, 10], pioneering work formally defined the task of panoptic symbol spotting [11] for floor plan CAD drawings and established the benchmark dataset FloorPlanCAD [11, 12, 13, 14, 15, 16]. However, the manual annotation of line-grained labels is a highly time-consuming and labor-intensive process, severely limiting the dataset’s scale and diversity. This bottleneck restricts the ability to train models that can generalize across diverse building types, varying spatial scales, complex layouts, and a wide range of architectural component categories, thereby hindering their applicability to real-world scenarios.
Figure 1: Illustration of Layer-Block Structure. (a) Layers: A CAD drawing is organized as multiple overlaid layers, each serving a distinct functional or categorical purpose. (b) Blocks: Frequently used elements are defined as reusable blocks, allowing for efficient replication and modification. Best viewed in color.
@@
Figure 2: An example of annotated drawings in ArchCAD-400K. Each geometric primitive is assigned both semantic and instance labels, with distinct semantic categories represented by different colors and instances visualized through convex hull masks. The chunks are extracted from complete drawings. Best viewed in color.
To address these limitations, we propose a highly efficient annotation pipeline for line-grained labels, which significantly reduces the annotation cost from 1,000 person-hours required for 16K data to 800 person-hours for 413K data. The core idea is to leverage the structured organization of floor plan drawings, including layers and blocks, to enable scalable and cost-effective large-scale annotation. Specifically, layers facilitate bulk annotation of specific categories through semantic aggregation (e.g., door and window layers), while blocks enable unified identification and annotation of repeated instances via an instance reuse mechanism. As shown in Figure 1, this layer-block structure, inherently designed to enforce drawing standards and ensure consistency in architectural documentation, forms the foundational framework for our automated annotation pipeline. In addition, for high data quality, we restrict data sources to completed drawings from leading design institutions and employ a fully vectorized annotation workflow to eliminate errors arising from image-to-vector conversions, with human experts reviewing and correcting automated results.
@@ 2.1 Floor Plan Datasets
Several datasets have been developed for floor plan analysis. SESYD [9] comprises 1,000 synthetic vectorized documents with ground truth annotations. FPLAN-POLY [10] contains 42 floor plans derived from images [17] for spatial relationship analysis. Cubicasa [3] provides 5,000 floor plan images annotated with over 80 object categories, focusing on residential layouts. RFP [18] includes 2,000 annotated floor plans with detailed room-level information for Asian residential buildings. FloorPlanCAD [11] offers 16,103 vector-graphic floor plans in “.svg" format, annotated across 35 object categories. LS-CAD [16] introduces a test set of 50 full-size CAD drawings with an average area of 1,000 square meters. However, these datasets are limited in scale and rely on labor-intensive annotation, hindering large-scale training.
@@ 2.2 Large-scale Vision Datasets
Large-scale vision datasets have significantly advanced deep learning models for complex visual understanding tasks[19, 20, 21, 22, 23, 24, 25, 26, 27], including floor plan recognition. Image-based datasets such as ImageNet [19] (14M images, 21K categories) and COCO [20] (300K images with object detection, segmentation, and captioning labels) have been instrumental in developing state-of-the-art models for object recognition, classification, and scene understanding. These datasets provide rich annotations that enhance model generalization, improving performance across general and domain-specific tasks. Nevertheless, large-scale vector-based datasets and efficient annotation processes for floor plan understanding remain scarce.
@@ 2.3 Panoptic Symbol Spotting
The panoptic symbol spotting task, initially proposed in [11], involves the simultaneous detection and classification of architectural symbols (e.g., doors, windows, stairs) in floor plan CAD drawings. While traditional methods [7] focus on countable instances (e.g., windows, tabls), Fan et al. [11] extended this to uncountable objects (e.g., walls, railings), inspired by [28]. They introduced PanCADNet, which integrates Faster R-CNN [29] for countable instances and Graph Convolutional Networks [30] for uncountable elements. Subsequently, Fan et al. [12] proposed CADTransformer, utilizing HRNetV2-W48 [31] and Vision Transformers [32] for primitive tokenization and embedding aggregation. Zheng et al. [15] adopted graph-based representations with Graph Attention Networks for semantic and instance-level predictions. Liu et al. [14] introduced SymPoint, exploring point set representations, later enhanced by SymPointV2 [13] through layer feature encoding and position-guided training. In contrast, CADSpotting [16] densely samples points along primitives and employs Sliding Window Aggregation for efficient panoptic segmentation of large-scale CAD drawings. While these methods perform well on FloorPlanCAD, they struggle on our more diverse, complex, and large-scale dataset, highlighting the need for robust and scalable solutions.
Figure 3: Overall pipeline of the annotation process. Primitives are labeled by automated methods followed by manual modification. Best viewed in color.
Figure 4: Annotation format of the ArchCAD-400K
@@ 3 Annotation pipeline of ArchCAD-400K
We carefully gathered over 11917 complete CAD drawings from the industry, covering a wide range of building types. From the initial selected drawings, 5,538 drawings were validated as strictly conforming to the layer-block organizational standards. Through automated annotation combined with expert refinement, we generated approximately 413,062 chunks with line-grained annotations, each sized at 14m×14m, aligning with the design of FloorPlanCAD [11]. The overall pipeline of the annotation process is shown in Figure 3. In the following, we outline the annotation process of ArchCAD-400K in depth.
@@ 3.1 Annotation Format
The ArchCAD-400K employs a structured annotation format tailored for the panoptic symbol spotting task, as defined in [11]. Each graphical primitive (e.g., line, arc, circle) in the drawing is characterized by a dual-identifier pair, $(l_{k},z_{k})$, where $l_{k}$ denotes its semantic category and $z_{k}$ represents the instance identifier. Primitives sharing the same $z_{k}$ value are considered to be part of the same instance. Figure 4 illustrates the annotation format for architectural layouts. This format uses an SVG file structure, representing elements such as paths and circles. Each graphical primitive is associated with a semantic label, an instance identifier, and an RGB color for visualization.
@@ 3.2 Layer-Block Standardization Screening
However, the feasibility of automated annotation depends on standardized drawings. Non-standard drawings can have ambiguous layer names and mixed primitives, leading to semantic confusion. Notably, professional design institutes typically maintain internal layering standards, which can be normalized and processed to establish machine-interpretable structures. Thus, we restrict our data to completed drawings from leading design institutions and apply a layer naming validation algorithm against a reference table, discarding drawings with over 5% deviation. This dual quality control ensures standardization of the input data, providing a solid foundation for future automated annotation based on layer-block standardization. The reference table can be found in our supplementary materials.
@@ 3.3 Automated Annotation with Expert Refinement
The automated annotation process, based on the layer-block standardization, significantly reduces the workload of manual annotation, providing an efficient foundation for large-scale drawing annotation. However, despite the substantial efficiency gains, the accuracy of automated annotation remains limited, particularly when handling non-standard drawings, which inevitably persist despite the standardization screening. Semantic ambiguities and mixed graphic primitives within layers may lead to annotation errors, making human intervention essential to ensure accuracy.
To address this, we engaged 10 professional architectural drafters with over 3 years of experience to refine the automatically generated annotations using an intuitive vector-based interface. Compared to the raster-based annotation method [11], such as polygon or bounding box labeling, directly operating on vector primitives not only eliminates potential errors introduced by raster-to-vector conversions but also overcomes the challenge of distinguishing overlapping instances in raster images, thereby reducing potential ambiguities. We created tools that automatically check CAD drawings against layer-block rules. Errors are highlighted directly on the drawings, allowing human corrections within seconds per file.
To achieve line-grained panoramic annotation, we have adopted a two-stage annotation system, as shown in Figure 3. In the semantic annotation stage, the system first classifies the layers based on their names and content into predefined semantic categories and displays the categories of each graphic primitive in professional CAD drawing software, allowing experts to directly correct inaccuracies through a vector-based operation interface. In the instance annotation stage, specific semantic categories are extracted and processed individually, isolating instances of different categories to reduce interference from surrounding information. The system automatically generates instance mask proposals based on existing block information in the drawings, which are then reviewed by experts who manually correct any errors. Through this two-stage annotation system, we completed the panoramic annotation of 5,538 drawings in approximately 800 hours, achieving an efficiency more than 10 times that of the FloorPlanCAD dataset. This process not only significantly enhances annotation efficiency but also ensures high accuracy and reliability.
@@ 4 Exploring ArchCAD-400K
FPLAN-POLY [10] Internet 48 <$5\times 10^{4}$ m2 Instance Human
SESYD [9] Synthetic 1000 <$5\times 10^{5}$ m2 Instance Human
FloorPlanCAD [11] Industry 16K $1.6\times 10^{6}$ m2 Panoptic Human
@@
We compare our ArchCAD-400K with existing datasets, as summarized in Table 1. Publicly available datasets, such as SESYD [9], FPLAN-POLY [10], and FloorPlanCAD [11], exhibit limitations in terms of source, scale, data format, annotation type and method, as well as the range of elements included in floor plans. These constraints affect their applicability to real-world architectural analysis.
@@ 4.1 Extended Building Diversity
Figure 5: (a) Different project types of drawings; (b) Area distribution of drawings.
Figure 6: Number of annotated primitives for 27 semantic classes. Bar colors match each class’s visualization.
@@ 4.2 Larger Data and Spatial Scale
ArchCAD-400K stands out for its unprecedented scale, comprising 5,538 complete drawings and 413,062 chunks, surpassing FloorPlanCAD’s 15,663 chunks by over 26 times. This substantial scale significantly enhances the robustness and generalization capabilities of AI models for floor plan analysis.
@@ 4.3 Diverse Element Types
Traditional floor plan datasets typically focus on common non-structural elements, such as furniture, equipment, and decorative features. While such categorizations are sufficient for residential interior layouts, they often fail to capture critical components when applied to more diverse building types, particularly in complex architectural or industrial settings. To address this limitation, our ArchCAD-400K adopts a comprehensive semantic categorization that extends beyond traditional architectural elements. Specifically, we introduce annotations for structural components (e.g., columns, beams, and holes) and drawing notations (e.g., axis lines, labels, and markers) in addition to non-structural elements. Structural components are fundamental to reconstructing the overall building structure as they define the load-bearing framework and spatial organization of the architectural design. In contrast, drawing notations are ubiquitous in real-world drawings and play a critical role in ensuring accurate interpretation and practical implementation. This tripartite categorization, encompassing non-structural, structural, and notation elements, enables a broader range of applications, including indoor navigation, architectural design, and structural analysis.
ArchCAD-400K provides detailed annotations for 27 categories, with the distribution of primitives across these categories visually represented in Figure 6. Notably, 14 of these categories each encompass more than 1 million primitives, underscoring the extensive scale of our dataset. Additionally, our dataset includes 7 countable instance categories. While this count is fewer than that of FloorPlanCAD, our dataset surpasses it in terms of diversity and complexity within each category, presenting a more formidable challenge. For instance, we consolidate various fine-grained furniture types from FloorPlanCAD, such as chair, table, and bed, into a single unified “furniture" category. Visual examples of the categories can be found in our supplementary materials.
@@ 5 Method
(a) DPSS Framework
(b) Adaptive fusion
Figure 7: The overall architecture of the proposed Dual-Pathway Symbol Spotter (DPSS). CAD drawings are encoded through both point backbone and image backbone. Features from dual pathways are aligned and fused in adaptive fusion module, then pass through a transformer decoder to generate mask predictions and classifications at primitive level. Best viewed in color.
To advance panoptic symbol spotting on large-scale datasets, we propose Dual-Pathway Symbol Spotter (DPSS), a novel framework designed for enhanced robustness and generalization. Unlike previous methods that rely on prior knowledge, such as color and layer information, DPSS eliminates these dependencies, improving adaptability across diverse datasets. As illustrated in Figure 7, the approach incorporates a dual-path feature extraction module, consisting of an image branch and a point cloud branch. The image branch encodes rendered vector graphics to capture global contextual features, while the point cloud branch represents primitives as points to extract fine-grained geometric details. An adaptive fusion mechanism aligns and integrates these multimodal features, ensuring effective feature representation for each primitive. Finally, a transformer decoder performs panoptic segmentation of the entire graphic, enabling precise symbol spotting in complex drawings.

### v3

# ArchCAD-400k: A Large-Scale CAD drawings Dataset and New Baseline for Panoptic Symbol Spotting

@@ Abstract
Recognizing symbols in architectural CAD drawings is critical for various advanced engineering applications. In this paper, we propose a novel CAD data annotation engine that leverages intrinsic attributes from systematically archived CAD drawings to automatically generate high-quality annotations, thus significantly reducing manual labeling efforts. Utilizing this engine, we construct ArchCAD-400k, a large-scale CAD dataset consisting of 413,062 chunks from 5538 standardized drawings, making it over 26 times larger than the largest existing CAD dataset. ArchCAD-400k boasts an extended drawing diversity and broader categories, offering line-grained annotations. Furthermore, we present a new baseline model for panoptic symbol spotting, termed Dual-Pathway Symbol Spotter (DPSS). It incorporates an adaptive fusion module to enhance primitive features with complementary image features, achieving state-of-the-art performance and enhanced robustness. Extensive experiments validate the effectiveness of DPSS, demonstrating the value of ArchCAD-400k and its potential to drive innovation in architectural design and construction.
@@ 1 Introduction
For a long time, CAD drawings have served as the universal language of architectural design, enabling seamless communication among designers, engineers, and construction personnel through standardized graphical primitives such as arcs, circles, and polylines. Accurate perception of these primitives in 2D CAD drawings is essential for various downstream applications, including automated drawing review and 3D building information modeling (BIM) kratochvila2024multi; park20213dplannet; kalervo2019cubicasa5k; liu2017raster, as this ability enables the optimization of CAD-based workflows while enhancing efficiency and precision in architectural design and construction processeszeng2019deep; yang2023vectorfloorseg.
Building upon earlier research in symbol spotting rezvanifar2019symbol; rezvanifar2020symbol; delalandre2010generation; rusinol2010relational, pioneering work formally defined the task of panoptic symbol spotting fan2021floorplancad for floor plan CAD drawings and established the benchmark dataset FloorPlanCAD fan2021floorplancad; fan2022cadtransformer; liu2024sympoint; liu2024symbol; zheng2022gat; mu2024cadspotting. However, the manual annotation of line-grained labels is a highly time-consuming and labor-intensive process, severely limiting the dataset’s scale and diversity. This bottleneck restricts the ability to train models that can generalize across diverse building types, varying spatial scales, complex layouts, and a wide range of architectural component categories, thereby hindering their applicability to real-world scenarios.
To address these limitations, we propose an efficient annotation pipeline for line-grained labels, reducing annotation cost from 1,000 person-hours for 16K data to 800 hours for 413K. The core idea is to leverage the structured organization of floor plan drawings, including layers and blocks, to enable scalable and cost-effective large-scale annotation. Layers support bulk labeling via semantic grouping (e.g., doors, windows), while blocks allow instance reuse for repeated elements. As shown in Figure 2, the layer-block structure, inherently designed to enforce drawing standards and ensure consistency in architectural documentation, forms the foundational framework for our automated annotation pipeline. To ensure high data quality, we use completed drawings from top design institutions and a fully vectorized annotation workflow, with expert review of automated outputs.
@@
Figure 1: An example of annotated drawings in ArchCAD-400k. Each geometric primitive is assigned both semantic and instance labels, with distinct semantic categories represented by different colors and instances visualized through convex hull masks. The chunks are extracted from complete drawings.
Figure 2: Layer-Block Structure: (a) Layers organize CAD drawings into functional categories. (b) Blocks define reusable elements for efficiency.
@@ 2.1 Floor Plan Datasets
Several datasets have been developed for floor plan analysis. SESYD delalandre2010generation comprises 1,000 synthetic vectorized documents with ground truth annotations. FPLAN-POLY rusinol2010relational contains 42 floor plans derived from images hilaire2006robust for spatial relationship analysis. Cubicasa kalervo2019cubicasa5k provides 5,000 floor plan images annotated with over 80 object categories, focusing on residential layouts. RFP lv2021residential includes 2,000 annotated floor plans with detailed room-level information for Asian residential buildings. FloorPlanCAD fan2021floorplancad offers 16,103 vector-graphic floor plans in “.svg" format, annotated across 35 object categories. LS-CAD mu2024cadspotting introduces a test set of 50 full-size CAD drawings with an average area of 1,000 square meters. However, these datasets are limited in scale and rely on labor-intensive annotation, hindering large-scale training.
@@ 2.2 Large-scale Vision Datasets
Large-scale vision datasets have significantly advanced deep learning models for complex visual understanding tasksdeng2009imagenet; lin2014microsoft; zhang2023movqa; kuznetsova2020open; zhou2017places; abu2016youtube; schuhmann2021laion; gupta2019lvis; liao2022kitti, including floor plan recognition. Image-based datasets such as ImageNet deng2009imagenet (14M images, 21K categories) and COCO lin2014microsoft (300K images with object detection, segmentation, and captioning labels) have been instrumental in developing state-of-the-art models for object recognition, classification, and scene understanding. These datasets provide rich annotations that enhance model generalization, improving performance across general and domain-specific tasks. Nevertheless, large-scale vector-based datasets and efficient annotation processes for floor plan understanding remain scarce.
@@ 2.3 Panoptic Symbol Spotting
The panoptic symbol spotting task, initially proposed in fan2021floorplancad, involves the simultaneous detection and classification of architectural symbols (e.g., doors, windows, stairs) in floor plan CAD drawings. While traditional methods rezvanifar2019symbol focus on countable instances (e.g., windows, tables), Fan et al. fan2021floorplancad extended this to uncountable objects (e.g., walls, railings), inspired by kirillov2019panoptic. They introduced PanCADNet, which integrates Faster R-CNN ren2016faster for countable instances and Graph Convolutional Networks kipf2016semi for uncountable elements. Subsequently, Fan et al. fan2022cadtransformer proposed CADTransformer, utilizing HRNetV2-W48 sun2019deep and Vision Transformers dosovitskiy2020image for primitive tokenization and embedding aggregation. Zheng et al. zheng2022gat adopted graph-based representations with Graph Attention Networks for semantic and instance-level predictions. Liu et al. liu2024symbol introduced SymPoint, exploring point set representations, later enhanced by SymPointV2 liu2024sympoint through layer feature encoding and position-guided training. In contrast, CADSpotting mu2024cadspotting densely samples points along primitives and employs Sliding Window Aggregation for efficient panoptic segmentation of large-scale CAD drawings. While these methods perform well on FloorPlanCAD, they struggle on our more diverse, complex, and large-scale dataset, highlighting the need for robust and scalable solutions.
Figure 3: Overall pipeline of the annotation process. Primitives are labeled by automated methods followed by manual modification.
@@ 3 Annotation pipeline of ArchCAD-400k
We carefully gathered over 11917 complete CAD drawings from the industry, covering a wide range of building types. From the initial drawings, 5,538 drawings were validated as strictly conforming to the layer-block organizational standards. Through automated annotation combined with expert refinement, we generated approximately 413,062 chunks with line-grained annotations, each sized at 14m×14m, aligning with the design of FloorPlanCAD fan2021floorplancad. The overall annotation pipeline of ArchCAD-400k is shown in Figure 3 and detailed below.
@@ 3.1 Annotation Format
The ArchCAD-400k employs a structured annotation format tailored for the panoptic symbol spotting task, as defined in fan2021floorplancad. Each graphical primitive (e.g., line, arc, circle) in the drawing is characterized by a dual-identifier pair, $(l_{k},z_{k})$, where $l_{k}$ denotes its semantic category and $z_{k}$ represents the instance identifier. Primitives sharing the same $z_{k}$ value are considered to be part of the same instance.
@@ 3.2 Layer-Block Standardization Screening
However, the feasibility of automated annotation depends on the standardization level of the drawings. Non-standard drawings can have ambiguous layer names and mixed primitives, leading to semantic confusion. Notably, professional design institutes typically maintain internal layering standards, which can be normalized and processed to establish machine-interpretable structures. Thus, we restrict our data to completed drawings from leading design institutions and apply a layer naming validation algorithm against a reference table, discarding drawings with over 5% deviation. This dual quality control ensures standardization of the input data, providing a solid foundation for future automated annotation based on layer-block standardization. The reference table is provided in Appendix A.
@@ 3.3 Automated Annotation with Expert Refinement
The automated annotation process based on layer-block standardization greatly reduces manual workload and enables large-scale annotation. However, its accuracy is limited, especially with non-standard drawings that escape standardization. Semantic ambiguities and mixed primitives within layers often cause errors, necessitating human correction. To improve accuracy, we engaged 10 experienced architectural drafters with over 3 years of experience to refine annotations using a vector-based interface. Unlike typical image-based annotation methods such as bounding boxes or polygons, direct vector editing avoids raster-to-vector conversion errors and resolves issues like overlapping instances.
We adopted a two-stage annotation system as shown in Figure 3. In the semantic stage, layers are categorized by names and content, with their semantics displayed in for expert review and correction. In the instance stage, key categories are isolated and instance masks are proposed from block information, then manually refined. We also developed tools to automatically check layer-block compliance, flagging errors for quick human correction. Using this system, we annotated 5,538 drawings in 800 hours, which is over 10 times more efficient than the FloorPlanCAD dataset while maintaining high accuracy and reliability.
Ethical and Copyright Considerations.
The dataset used in this work is derived from architectural or design drawings, which may contain potentially sensitive or proprietary information. To ensure compliance with ethical and copyright standards, we apply strict data anonymization before training or release, including removal of identifiable text and irreversible obfuscation of metadata, so that the original content of any individual project cannot be reconstructed. All data is used solely for research under academic fair use, and no raw data that may infringe copyright or IP rights is released. These measures ensure compliance with ethical standards and protection of PII.
@@ 4 Exploring ArchCAD-400k
FPLAN-POLY rusinol2010relational Internet 48 $5\times 10^{4}$ m2 Instance Human
SESYD delalandre2010generation Synthetic 1000 $5\times 10^{5}$ m2 Instance Human
FloorPlanCAD fan2021floorplancad Industry 16K $1.6\times 10^{6}$ m2 Panoptic Human
@@
We compare our ArchCAD-400k with existing datasets, as summarized in Table 1. Publicly available datasets, such as SESYD delalandre2010generation, FPLAN-POLY rusinol2010relational, and FloorPlanCAD fan2021floorplancad, exhibit limitations in terms of source, scale, data format, annotation type and method, as well as the range of elements included in floor plans. These constraints affect their applicability to real-world architectural analysis.
@@ 4.1 Extended Building Diversity
Figure 4: (a) Different project types of drawings; (b) Area distribution of drawings; (c) Number of annotated primitives for 27 semantic classes, where bar colors match each class’s visualization.
@@ 4.2 Larger Data and Spatial Scale
ArchCAD-400k stands out for its unprecedented scale, comprising 5,538 complete drawings and 413,062 chunks, surpassing FloorPlanCAD’s 15,663 chunks by over 26 times. This substantial scale significantly enhances the robustness and generalization capabilities of models for the spotting task.
@@ 4.3 Diverse Element Types
Traditional floor plan datasets typically focus on common non-structural elements, such as furniture, equipment, and decorative features. While such categorizations are sufficient for residential interior layouts, they often fail to capture critical components when applied to more diverse building types, particularly in complex architectural or industrial settings.
To address this limitation, our ArchCAD-400k adopts a comprehensive semantic categorization that extends beyond traditional architectural elements. Specifically, we introduce annotations for structural components (e.g., columns, beams, and holes) and drawing notations (e.g., axis lines, labels, and markers) in addition to non-structural elements. Structural components are fundamental to reconstructing the overall building structure as they define the load-bearing framework and spatial organization of the architectural design. In contrast, drawing notations are ubiquitous in real-world drawings and play a critical role in ensuring accurate interpretation and practical implementation. This tripartite categorization, encompassing non-structural, structural, and notation elements, enables a broader range of applications, including indoor navigation, architectural design, and structural analysis.
ArchCAD-400k provides detailed annotations for 27 categories, with the distribution of primitives across these categories visually represented in Figure 4. Notably, 14 of these categories each encompass more than 1 million primitives, underscoring the extensive scale of our dataset. Additionally, our dataset includes 7 countable instance categories. While this count is fewer than that of FloorPlanCAD, our dataset surpasses it in terms of diversity and complexity within each category, presenting a more formidable challenge. For instance, we consolidate various fine-grained furniture types from FloorPlanCAD, such as chair, table, and bed, into a single unified “furniture" category. Visual examples of the categories can be found in Appendix A.
@@ 5 Method
To enhance panoptic symbol spotting on large-scale datasets, we propose Dual-Pathway Symbol Spotter (DPSS), a robust framework free from prior knowledge like color or layer cues. As shown in Figure 5, DPSS employs a dual-path feature extractor: an image branch encodes rendered graphics for global context, while a point cloud branch captures geometric details by treating primitives as points. An adaptive fusion module aligns these multimodal features, and a transformer decoder performs panoptic segmentation for precise symbol spotting in complex diagrams.